# Counterexamples to two conjectures on Venn diagrams

### v3

@@ 3.1.2. Step (ii): Computing all $C$-quadrangulations
For a fixed cycle $C$ in $Q_{n-1}$ as computed in the previous step, we aim to compute all $C$-quadrangulations. This is done recursively, by filling the inside of $C$ with 4-faces, one at a time. In each step, the current partial $C$-quadrangulation $P$ has an inner face $B$ that still needs to be quadrangulated; see Figure 8 (b). For this we consider every edge $e$ along $B$, and every 4-cycle $F$ in $Q_{n-1}$ that contains $e$ (there are $n-3$ such 4-cycles). Every such 4-cycle $F$ containing $e$ has two additional vertices $u$ and $v$. We locate $u$ and $v$ in $P$ and distinguish the following cases: If $u$ and $v$ are on $B$ but in the wrong cyclic order, then $F$ is excluded, i.e., it will not be added to $P$ as a 4-face. Similarly, if $u$ or $v$ are present in $P$ but not on $B$, then $F$ is excluded. These two cases are shown in Figure 9 (c), respectively. If $u$ and $v$ are on $B$ but not next to $e$ (Figure 9 (b)), then $F$ is ignored for this recursion step (because including $F$ would split the inner face bounded by $B$ into two separate faces). If $u$ and $v$ are next to $e$ on $B$ or not in $P$ (Figure 9 (a) shows the six possible such cases), then $F$ is eligible for a branching step. Specifically, we branch and recurse on the two possibilities of either including $F$, updating $B$ accordingly, or excluding $F$, which does not change $B$. Some care has to be taken to choose a clever branching order to avoid recomputing the same $C$-quadrangulation along different insertion orderings of its 4-faces, and also to keep the recursion tree small, i.e., to enforce and detect dead ends quickly. Such a dead end would be an edge $e$ on $B$ for which all 4-cycles $F$ containing $e$ are excluded (apart possibly from one that can be included ‘behind’ $e$).
@@ 4. Proof of Theorem 6
Proof of Theorem 6.
We inductively construct a family of non-simple monotone $n$-Venn diagrams $D_{n}$ for $n\geq 2$; see Figures 10 and 11. These are not the counterexamples yet, but those will be built from $D_{n}$ later by small modifications. Specifically, we describe $D_{n}$ as a wire diagram, i.e., each crossing occurs within a horizontally separated narrow vertical strip, and outside of these strips the $n$ curves are aligned to horizontal lines at heights $1,2,\ldots,n$. Furthermore, the diagram wraps around cylindrically at the left and right boundary. The curves are labeled $1,2,\ldots,n$ from bottom to top at the boundaries, according to their height.
For the induction basis $n=2$, we take $D_{2}$ as the wire diagram obtained from the unique Venn diagram with 2 curves, which has 2 crossings; see Figure 10. For the induction step $n\rightarrow n+1$, let $D_{n}^{-}$ be the diagram obtained from $D_{n}$ by removing the rightmost crossing. To construct $D_{n+1}$, we take a copy of $D_{n}^{-}$ vertically aligned at heights $1,\ldots,n$, a second copy of $D_{n}^{-}$ but reversed (mirrored along a vertical line, i.e., left and right are interchanged) to the right of the first copy and vertically aligned at heights $12,\ldots,n+$ 1; see Figure 11. We then add the curve $n+1$ starting on the left at height $n+1$ above the first copy of $D_{n}^{-}$, staying there until its right side, then crossing below the reversed second copy of $D_{n}^{-}$ to height 1, staying there until its right side, and finally crossing back up to height $n+1$ at the right boundary. The curves $1,\ldots,n$ are connected between the two copies of $D_{n}^{-}$ through one additional crossing with the curve $n+1$, and to the right of the second copy through a second additional crossing with the curve $n+1$, creating two additional crossings in total, so that the ordering of curves on the right boundary becomes $11,\ldots,n+$ 1, as desired. One can easily verify the following properties inductively:
@@
From (ii) we see in particular that $D_{n}$ has a perfect matching and a Hamilton cycle, i.e., it is not a counterexample to Conjecture 5. However, for $n\geq 4$ we can modify $D_{n}$ to obtain a diagram $D_{n}^{*}$ as follows; see Figure 12: From the first crossing involving all $n$ curves, we slightly shift down the curve $n-1$, and from the second crossing involving all $n$ curves, we slightly shift up the curve $n-1$, so that each of these two crossings is split into two, one of degree 4 and the other of degree $22n-$ 2 (this is possible because of (iii)). By (i), the number of vertices of $D_{n}^{*}$ is $22^{n-1}+$ 2, an even number. Furthermore, in the graph of $D_{n}^{*}$, let $U$ be the independent set of all degree 4 vertices and let $\overline{U}$ be its complement. From (ii) we obtain that $|U|=2^{n-1}/2+2=2^{n-2}+2$ and $|\overline{U}|=2^{n-1}/2=2^{n-2}$, and consequently $|U|>|\overline{U}|$. This inequality and the fact that $U$ is an independent set shows that $D_{n}^{*}$ does not have a perfect matching. This is the easy direction of Tutte’s theorem. ∎
Figure 12. Modification of the diagram $D_{n}$ to obtain $D_{n}^{*}$ by slightly shifting the curve ${n-1}$.
@@ 5. Open questions
A natural question that remains open is whether Winkler’s conjecture holds for monotone Venn diagrams. We believe that the answer is ‘no’, but there are no counterexamples for $n\leq 6$ (recall Figure 4). One idea to find a monotone $n$-Venn quadrangulation without a perfect matching for $n\geq 7$ would be to extend the local obstruction $H$ highlighted in Figure 3, but so far we have not succeeded in this endeavor.
Computationally, it could be interesting to find all non-simple Venn diagrams for $n=4$ or $n=5$ curves and analyze their properties. What can be said about the asymptotic number of $n$-Venn diagrams as $n$ grows, or about the proportion of diagrams with certain properties?
In addition to these questions, there is a large number of other intriguing problems on Venn diagrams, concerning symmetry, convexity, number of crossings, restrictions on the allowed shapes of the curves (triangles, rectangles etc.), to be found in the survey [RW97]. For example, is there a simple $n$-Venn diagram with rotational symmetry made of $n$ congruent curves for every prime number $n$ (see [GKS04, KRSW04, MR14])?

### v4

@@ 3.1.2. Step (ii): Computing all $C$-quadrangulations
For a fixed cycle $C$ in $Q_{n-1}$ as computed in the previous step, we aim to compute all $C$-quadrangulations. This is done recursively, by filling the inside of $C$ with 4-faces, one at a time. In each step, the current partial $C$-quadrangulation $P$ has an inner face $B$ that still needs to be quadrangulated; see Figure 10 (b). For this we consider every edge $e$ along $B$, and every 4-cycle $F$ in $Q_{n-1}$ that contains $e$ (there are n3n-3 such 4-cycles). Every such 4-cycle $F$ containing $e$ has two additional vertices $u$ and $v$. We locate $u$ and $v$ in $P$ and distinguish the following cases: If $u$ and $v$ are on $B$ but in the wrong cyclic order, then $F$ is excluded, i.e., it will not be added to $P$ as a 4-face. Similarly, if $u$ or $v$ are present in $P$ but not on $B$, then $F$ is excluded. These two cases are shown in Figure 11 (c), respectively. If $u$ and $v$ are on $B$ but not next to $e$ (Figure 11 (b)), then $F$ is ignored for this recursion step (because including $F$ would split the inner face bounded by $B$ into two separate faces). If $u$ and $v$ are next to $e$ on $B$ or not in $P$ (Figure 11 (a) shows the six possible such cases), then $F$ is eligible for a branching step. Specifically, we branch and recurse on the two possibilities of either including $F$, updating $B$ accordingly, or excluding $F$, which does not change $B$. Some care has to be taken to choose a clever branching order to avoid recomputing the same $C$-quadrangulation along different insertion orderings of its 4-faces, and also to keep the recursion tree small, i.e., to enforce and detect dead ends quickly. Such a dead end would be an edge $e$ on $B$ for which all 4-cycles $F$ containing $e$ are excluded (apart possibly from one that can be included ‘behind’ $e$).
@@ 4. Proof of Theorem 7
Proof of Theorem 7.
We inductively construct a family of non-simple monotone $n$-Venn diagrams $D_{n}$ for $n\geq 2$; see Figures 12 and 13. These are not the counterexamples yet, but those will be built from $D_{n}$ later by small modifications. Specifically, we describe $D_{n}$ as a wire diagram. The curves are labeled $1,2,\ldots,n$ from bottom to top at the left and right boundary, according to their height.
@@
Figure 14. Modification of the diagram $D_{n}$ to obtain $D_{n}^{*}$ by slightly shifting the curve n1{n-1}.
For the induction basis $n=2$, we take $D_{2}$ as the wire diagram obtained from the unique Venn diagram with 2 curves, which has 2 crossings; see Figure 12. For the induction step $n\rightarrow n+1$, let $D_{n}^{-}$ be the diagram obtained from $D_{n}$ by removing the rightmost crossing. To construct $D_{n+1}$, we take a copy of $D_{n}^{-}$ vertically aligned at heights $1,\ldots,n$, a second copy of $D_{n}^{-}$ but reversed (mirrored along a vertical line, i.e., left and right are interchanged) to the right of the first copy and vertically aligned at heights $2,\ldots,n+1$; see Figure 13. We then add the curve $n+1$ starting on the left at height $n+1$ above the first copy of $D_{n}^{-}$, staying there until its right side, then crossing below the reversed second copy of $D_{n}^{-}$ to height 1, staying there until its right side, and finally crossing back up to height $n+1$ at the right boundary. The curves $1,\ldots,n$ are connected between the two copies of $D_{n}^{-}$ through one additional crossing with the curve $n+1$, and to the right of the second copy through a second additional crossing with the curve $n+1$, creating two additional crossings in total, so that the ordering of curves on the right boundary becomes $1,\ldots,n+1$, as desired. One can easily verify the following properties inductively:
@@
From (ii) we see in particular that $D_{n}$ has a perfect matching and a Hamilton cycle, i.e., it is not a counterexample to Conjecture 6. However, for $n\geq 4$ we can modify $D_{n}$ to obtain a diagram $D_{n}^{*}$ as follows; see Figure 14: From the first crossing involving all $n$ curves, we slightly shift down the curve n1n-1, and from the second crossing involving all $n$ curves, we slightly shift up the curve n1n-1, so that each of these two crossings is split into two, one of degree 4 and the other of degree 2n22n-2 (this is possible because of (iii)). By (i), the number of vertices of $D_{n}^{*}$ is $2^{n-1}+2$, an even number. Furthermore, in the graph of $D_{n}^{*}$, let $U$ be the independent set of all degree 4 vertices and let $\overline{U}$ be its complement. From (ii) we obtain that $|U|=2^{n-1}/2+2=2^{n-2}+2$ and $|\overline{U}|=2^{n-1}/2=2^{n-2}$, and consequently $|U|>|\overline{U}|$. This inequality and the fact that $U$ is an independent set shows that $D_{n}^{*}$ does not have a perfect matching. This is the easy direction of Tutte’s theorem. ∎
@@ 5. Open questions
Computationally, it could be interesting to find all non-simple Venn diagrams for $n=4$ or $n=5$ curves and analyze their properties.
Also, what can be said about the asymptotic number of $n$-Venn diagrams as $n$ grows, or about the proportion of diagrams with certain properties (recall Figure 6)?
There is a large number of other intriguing problems on Venn diagrams, concerning symmetry, convexity, number of crossings, restrictions on the allowed shapes of the curves (triangles, rectangles etc.), to be found in the survey [RW97]. For example, is there a simple $n$-Venn diagram with rotational symmetry made of $n$ congruent curves for every prime number $n$ (see [GKS04, KRS+04, MR14])?